OpenStudy (anonymous):

help meee! find the volume of a spehere with a radius of 9cm. dont gie answer show me how to do it. please(:

OpenStudy (nurali):

volume of a sphere= 4/3 * pi * r^3 pi=3.14 here r = 9units so volume of a sphere= 4/3 * 3.14 * (9)^3 volume of a sphere= 4/3 * 3.14 * 729 volume of a sphere = 3052.08 cubic units

OpenStudy (nathan917):

volume of a sphere= 4/3 * pi * r^3 pi=3.14 here r = 9units so volume of a sphere= 4/3 * 3.14 * (9)^3 volume of a sphere= 4/3 * 3.14 * 729 volume of a sphere = **** she said she did not what the answer man.

OpenStudy (nathan917):

She wanted to solve it by herself.